(b) For the given sequence 1/n, 1/2n, 1/4n,… Find the 16th term. Find sum of first sixteen terms where n is your arid number i.e. 19-arid-234 take n=234


1
Expert's answer
2021-07-16T02:53:29-0400

(b)


"\dfrac{1}{n},\dfrac{1}{2n},\dfrac{1}{4n},\dfrac{1}{8n}...."

This is GP

So,


First term is "a=\dfrac{1}{n}"


and common ratio "r=\dfrac{1}{2}"


"p^{th}" term of GP

"a_p=ar^{p-1}"


So, 16th term

"a_{16}=ar^{16-1}=ar^{15}=\dfrac{1}{n}\cdot (\dfrac{1}{2})^{15}=\dfrac{1}{32768n}"

and for n= 234

"a_{16}=\dfrac{1}{32768\times 234}=\dfrac{1}{7667712}"



and sum of p terms of GP

"S_p=\dfrac{a(1-r^p)}{1-r}"


So, sum of 16 terms of GP is

"S_{16}=\dfrac{\frac{1}{n}(1-(\frac{1}{2})^{16})}{1-\frac{1}{2}}=\dfrac{2}{n}(1-(\frac{1}{2})^{16})"

for n= 234

"S_{16}=\dfrac{2}{234}[1-\dfrac{1}{65536}]=\dfrac{21845}{2555904}"